OpenStudy (ashleneperalez):

Determine which ordered pair is a solution of Y=-5x+10 (-15,5) (2,0) (-1,-5) (3,25)

rebeccaxhawaii (rebeccaxhawaii):

plug them in

OpenStudy (anonymous):

(x,y) plug them in their corresponding slots and solve.

OpenStudy (anonymous):

(-1,-5) for example. y=-5x+10 -5=-5(-1)+10 -5=5+10 -5≠10

OpenStudy (anonymous):

so we know it can't be (-1,-5) try it with the rest and find the answer easily

OpenStudy (anonymous):